OpenStudy (anonymous):

i am trying to find out how much i need to deposit to get 25000 in 6years 5% compounded semiannually, could you please help me with the correct formula? i have tried but to no avail thank you in advance

OpenStudy (anonymous):

As I said in your previous question, the formula is Effective interest rate r = (1 + APR/[# of times compounded])^[# of times compounded] - 1 = (1 + 0.05/2)^2 - 1 = 0.050625 or 5.0625% per annum. So over 6 years you start with an amount D, which becomes at the end of: 1 year D(1.050625) 2 year D(1.050625)^2 3 year D(1.050625)^3 4 year D(1.050625)^4 5 year D(1.050625)^5 6 year D(1.050625)^6 So you have an unknown variable D which is multiplied by (1.050625)^6 to get D + 25000. Then: D(1.050625)^6 = D + 25000 D(1.050625)^6 - D = 25000 1.34488882424*D - D = 25000 0.34488882424*D = 25000 D = 25000/0.34488882424 ~= $8622.22

OpenStudy (anonymous):

nicely explained rav, i tried your other way but was struggling with it, just thought i had worded it incorrectly

OpenStudy (anonymous):

Thanks, sorry I made a mistake at the very end, I multiplied instead of dividing... that number should be $72487.13 not $8622

OpenStudy (anonymous):

But the rest of it is correct don't worry

OpenStudy (anonymous):

lol